Circular No. 6213

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA CASSIOPEIAE 1995 K. Hirosawa, Inazawa, Aichi, reports the discovery by Minoru Ya-

mamoto, Okazaki, Aichi, of a possible nova on Aug. 24 at α = 1^{h}05^{m}.1, δ = +54^{o}02′ (equinox 2000.0). S. Nakano, Sumoto, communicates the following precise position, obtained on Aug. 26 by T. Kojima: α = 1^{h}05^{m}05^{s}.37, δ = +54^{o}00′40′′.5 (equinox 2000.0, GSC). Hirosawa also communicates positions obtained on Aug. 26 by M. Iida (end figures 05^{s}.66, 42′′.5) and by A. Sugie (end figures 05^{s}.35, 40′′.0). Magnitude estimates: July 24, [10 (Yamamoto); Aug. 24.57 UT, 9.2 (Yamamoto, 200-mm f/4.0 lens and PO0 filter, T-Max 400 film); 26.58, 9.0 (Iida, Nagano, 0.16-m reflector + CCD, no filter); 26.62, 9.3 (S. Takahashi, 0.15-m refractor, visual); 26.62, 9.2 (Kojima, Chiyoda, 0.25-m f/6 reflector + CCD, V filter); 26.65, 9.37 (Sugie, Dynic, 0.6-m reflector + CCD, no filter). G. V. Williams, Harvard-Smithsonian Center for Astrophysics, has identified the probable precursor on the Digital Sky Survey (epoch 1954.752, Palomar red plate, mag 18–19) at end figures 05^{s}.40, 40′′.6 (11 GSC reference stars, rms residual 0′′.3).

EDITORIAL NOTICE Users of these Circulars are reminded that the information contained

in them may be freely used, provided that any republication includes an appropriate acknowledgment, while the actual form of the Circulars and sections thereof is subject to the usual copyright restrictions. These remarks apply both to the printed and electronic versions. The indiscriminate reproduction of whole or partial Circulars in electronic bulletin boards and information services is inappropriate. Reproduction of this “form of expression” should be restricted to “fair use” by subscribers and their close associates, or by application to the undersigned for permission to reproduce all or part of a Circular under occasional, special circumstances. These Circulars are distributed by subscription. Issues more than one year old are available free of charge in the World Wide Web (http://cfa-

www.harvard.edu/cfa/ps/cbat.html) in both ASCII and .dvi form but

are subject to the same redissemination policy. Appropriate forms of acknowledgment for the use of information in a particular Circular can be to the Circular No. itself, to the author of the information extracted (information not otherwise attributed being considered as provided by the editor), or to both. The willful republication of cometary orbital elements (for example) without an indication of their source is a clear act of plagiarism.

1995 August 26 Brian G. Marsden